0 0
Add a comment Improve this question Transcribed image text
Answer #1

d = L/2


I = (1/3)*m*L^2


T = 2*pi*sqrt(I/m*g*d)


I / mgd = T^2/4pi^2 = (1.7*1.7)/(4*3.14*3.14) = 0.07327


(1/3)*m*L^2/ m*g*(L/2) = 0.07327


(2/3)*L/g = 0.07327

L = 1.077 m

-------------


w = 2*pi/T = (2*3.14)/1.7 = 3.69 rad/s

A = L*theta = 1.077*4.3*(3.14/180) = 0.081 m

KE max = 0.5*I*w^2 = 0.5*(1/3)*m*L^2*w^2

KE = 0.5*(1/3)*0.16*1.077*1.077*3.69*3.69

KE = 0.421 J